(X-2)/(x+1)-(3-12x)/(2x^2-x-3)
Answer by sachi(548)

as2x^2-x-3=2x^2+2x-3x-3=2x(x+1)-3(x+1)=(x+1)(2x-3)

now (X-2)/(x+1)-(3-12x)/(2x^2-x-3)=(x-2)/(x+1)-(3-12x)/(x+1)(2x-3)
=[(x-2)(2x-3)-(3-12x)]/(x+1)(2x-3)
=[2x^2-3x-4x+6-3+12x]/(x+1)(2x-3)
=[2x^2+5x+3]/(x+1)(2x-3)
=[2x^2+2x+3x+3]/(x+1)(2x-3)
=[2x(x+1)+3(x+1)]/(x+1)(2x-3)
=(x+1)(2x+3)/(x+1)(2x-3)
=(2x+3)/(2x-3) ans